# Systematic Study of Amorphous ABC Heterostructures at the Atomic Scale as a Nonlinear Optical Metamaterial

### v1

@@
Abstract
@@
The systematic exploration of ABC type heterostructures reveals that nanoscale morphological modification markedly improves nonlinear optical properties to maximize the artificial bulk second-order susceptibility. These amorphous birefringent heterostructures are fabricated through cyclic plasma-enhanced atomic layer deposition of three oxides, effectively breaking centrosymmetry. We observe a dependence of optical nonlinearity on the thickness variation of three constituent materials: SiO2 (A), TiO2 (B), and Al2O3 (C), ranging from tens of nanometers to the atomic scale, and these materials exhibit second-order susceptibility at their interfaces. Our findings reveal that the enhancement of nonlinear optical properties is strongly correlated with a high density of layers and superior interface quality, where the interface second-order nonlinearity transitions to bulk-like second-harmonic generation. An effective bulk second-order susceptibility of $\chi_{zzz}=2.0\pm 0.2$ pm/V is achieved, comparable to typical values for conventional monocrystalline nonlinear materials.
second harmonic generation, surface second-order susceptibility, interface second-order susceptibility, nonlinear metamaterial, ABC heterostructure, nanolaminates, quaternary oxide, tricolour heterostructure
I Introduction
@@
example2023The rise of atomic layer deposition (ALD) has revolutionized numerous technologies, including transistors ALD-tranzistor , memory devices ALD-memory , solar cells solarcell , catalysts catalyst , batteries batteries , conformal optical coatings AdrianaALD , and many others. However, its impact on the field of nonlinear optics remains largely untapped. Our research investigates the potential of ALD-deposited amorphous heterostructures, specifically ABC type nanolaminates consisting of three distinct dielectrics SiO2 (A), TiO2 (B), and Al2O3 (C), to create CMOS-compatible second-order nonlinear metamaterials. The core concept of metamaterials is to engineer materials with properties that are not found in nature by designing their structure at a scale smaller than the wavelength of light they interact with. Atomic layer deposition enables unparalleled precision in the fabrication of the structures, building them one atomic layer at a time with exceptional control and uniformity. Previous studies on layered composites with nonzero bulk second-order susceptibility assumed that only non-centrosymmetric constituents contribute to nonlinearity. However, studies neglected interface contributions to the nonlinear susceptibility of the heterostructure Boyd:94 . Our goal is to leverage the often overlooked interface, or surface, second-order nonlinearity heinz1991second ; shen1989optical by stacking multiple thin oxide layers in 2D heterostructures, thus significantly enhancing the effective nonlinear response ABC-Alloatti ; ABC-Clemmen . Impressively, these heterostructures exhibit a robust, bulk-like nonlinear response, even though all constituent oxides lack inherent bulk second-order nonlinearity and are amorphous, as shown in Fig. 1. This produces a highly distinctive material where second-order nonlinear processes arise not from bulk nonlinearity but from surface nonlinearity at the interfaces between centrosymmetric constituents. Previous studies on ternary oxide heterostructures (Al2O3/TiO2 and SiO2/HfO2) have demonstrated superior structural and optical properties of ALD-grown thin films, achieving precise control down to the atomic scale palabi ; PallabiHfO2 . These structures offer temperature and chemical stability, low-temperature growth requirements of $100\penalty 10000\ ^{\circ}\text{C}$, and conformal deposition on a wide range of nano- and micro-structured substrates, without the need to account for substrate compatibility or crystal growth orientation, unlike nonlinear III–V semiconductors such as GaAs GaAs . Moreover, they bypass the high-temperature manufacturing requirements of lithium niobate on insulator (LNOI), another widely used nonlinear platform LNOI . They also provide the flexibility to engineer the effective refractive index for specific applications by adjusting the ratio of constituent oxides palabi . ABC stacks exhibit outstanding optical transparency across the visible spectrum, surpassing the performance of typical III-V technologies. Moreover, the ALD deposition technique and the dielectrics used are standard in the semiconductor industry. Altogether, these advantages make ABC heterostructures a powerful platform for integrating photonic structures on a chip using only commonly available oxides and established semiconductor processes.
@@
Figure 1: The visualization of the ABC type heterostructures with an optical axis perpendicular to the layers. Changing the density of layers influences nonlinear optical properties.
@@
In bulk media with inversion symmetry, second-order nonlinear processes are forbidden under the electric dipole approximation, leading to zero bulk second-order susceptibility, $\chi_{\text{bulk}}^{(2)}=0$ Shen ; Boyd1992 , and no second harmonic enhancement is observed with varying thickness in pure samples of centrosymmetric materials A, B, or C. However, at the interface between two centrosymmetric media, the inversion symmetry, defined by the transformation $(x,y,z)\to(-x,-y,-z)$, is broken. This symmetry breaking results in a nonzero surface second-order susceptibility, $\chi_{\text{s}}^{(2)}\neq 0$, enabling electric-dipole-allowed second-harmonic generation (SHG) at surfaces and interfaces Shen1989 ; GuyotSionnest1987 . This principle is the key behind ABC type heterostructures, which consist of nanometer-thick amorphous (centrosymmetric) layers of three materials: A, B, and C. The inversion symmetry is broken locally at each AB, BC, and CA interface. In the case of AB stacks, the centrosymmetry is broken only locally, not globally. However, in ABC stacks, the inversion symmetry is broken both locally at the interfaces and globally, prohibiting total destructive interference of the second harmonic, resulting in a net $\chi_{\text{s}}^{\text{ABC}}$ for the ABC stack. The literature suggests that ABCD stacks, such as those incorporating SiO2/TiO2/Al2O3 with the addition of Hf2O2, may outperform ABC stacks. However, this potential enhancement remains experimentally unverified ABC-Alloatti ; ABC-Clemmen . Unfortunately, clear studies are lacking on which combination of inversion-symmetric materials offers the best nonlinear responses. A promising approach is to select materials with high dielectric contrast heinz1991second . However, experimental values for surface second-order susceptibility of materials and interfaces are scarce rodriguez2008calibration ; wang2010multipolar .
@@
Surface nonlinearity arises from three main physical origins. First, the electric dipole contribution due to symmetry breaking at interfaces. Second, a field discontinuity at the interface, where the normal component of the electric field ($E_{z}$) changes rapidly, contributes to the surface nonlinearity via the electric quadrupole (nonlocal) term. Third, structural disparities between materials, such as liquid/solid interfaces, contribute to bulk-like surface nonlinearity even when dielectric constants are matched. The second and third contributions are known as electric quadrupole or nonlocal contributions to surface nonlinear susceptibility shen1989optical . Electric dipoles dominate second-harmonic generation, although small contributions from electric quadrupoles and magnetic dipoles are not ruled out multipolar1 ; multipolar2 .
ABC type heterostructures have birefringent properties hermans2019 . For monocrystals, birefringence can be explained by the anisotropic electrical properties of the molecules that make up the crystals. However, birefringence can also be caused by anisotropy if the material is arranged on an order of magnitude larger than the size of the molecules and this distance is still smaller than the wavelength of light. The ABC type heterostructure, composed of three materials arranged in thin layers, acts as an effective medium exhibiting form birefringence and functioning as a negative uniaxial crystal. At the same time, the entire structure is amorphous, and the optical axis is perpendicular to the layers born2013principles ; Formbiro .
@@
In this work, we present experimental investigations of ABC heterostructures in the form of SiO2/TiO2/Al2O3, deposited via atomic layer deposition. We systematically vary the thickness of each constituent material to evaluate its impact on the efficiency of second-harmonic generation. We systematically explore the structural, linear, and nonlinear optical properties for each morphological composition.
II Deposition and Characterization
@@
Plasma-enhanced atomic layer deposition (PEALD) was performed using a SILAYO-ICP330 system (Sentech Instruments GmbH) to fabricate SiO2/TiO2/Al2O3 heterostructures. The deposition utilized oxygen plasma as the co-reactant with the following precursors: bis(diethylamino)silane (BDEAS) for SiO2, tetrakis(dimethylamino)titanium (TDMAT) for TiO2, and trimethylaluminum (TMA) for Al2O3. The inductively coupled plasma (ICP) source was operated at $100\penalty 10000\ \text{W}$ for all processes. The reactor chamber was kept at a temperature of $100\penalty 10000\ ^{\circ}\text{C}$. A 15-minute stabilization period was implemented after sample loading prior to deposition to ensure thermal stability and uniformity. Nitrogen (N2) served as the precursor carrier and purge gas. The growth per cycle (GPC) is determined for single-layer depositions of approximately 20-50 nm. All deposition parameters are summarized in Table 1.
@@
To examine the samples, high-resolution scanning transmission electron microscopy (HR-STEM) studies were carried out using a Themis Z(3.1) 80-300 TEM instrument (Thermo Fisher Scientific) with probe Cs-aberration correction, operating at 300 nm. The microscope is equipped with a Super-XTM (TFS) Energy Dispersive X-ray (EDX) detector and an a Gatan Image Filter (GIF) Continuum 1065ER used for Electron Energy Loss Spectroscopy (EELS). The samples were prepared by Focused Ion Milling (FIB) lift-out technique using a Helios G4 UC FIB-SEM device operated at 30, 16 and 5 kV.
@@
The refractive indices were determined for each structure using an SE850 DUV variable angle spectroscopic ellipsometer (Sentech Instruments GmbH), and the data were evaluated in the wavelength range from 400nm400\ $\text${nm} to 1040nm1040\ $\text${nm} with a uniaxial Sellmeier dispersion model.
@@
The following optical setup was used to investigate the nonlinear optical response of the ABC layers. The femtosecond laser (PHAROS-SP, Light Conversion) emits pulses with full-width half-maxima (FWHM) of $177\text{\,}\mathrm{fs}$ with a central wavelength of 1032 nm and with an average power of up to $1.5\text{\,}\mathrm{W}$ at a repetition rate of $200\text{\,}\mathrm{kHz}$. Polarization-based measurements were performed by a femtosecond Satsuma laser (Amplitude) with full-width half maxima of $270\text{\,}\mathrm{fs}$, wavelength of 1031.4 nm and the repetition rate of $1\text{\,}\mathrm{MHz}$, Fig. 2. Power and polarization are controlled by a half-wave plate HW1 (WPH05M-1030, Thorlabs), polarizer P1 (VA5, Thorlabs), and half-wave plate HW2 (Thorlabs AHWP05M-950). The laser beam is focused on the sample by the plano-convex lens L1 (LA1172-B-ML, Thorlabs) with a focal length of 400 mm, producing a beam waist of $w_{0}=55\ \upmu\text{m}$. This leads to the confocal parameter of 18, that is, $2\timesRayleigh$ length. The loosely focused beam enables the front and back sides of the sample to remain in focus during the rotation process; the sample is 1 mm thick. A long pass filter LP (FELH0900, Thorlabs) is positioned behind the lens L1 to remove any parasitic light, especially any SHG created at other optical elements. The sample is mounted onto an automated rotation stage, and its position is precisely aligned so that the laser beam stays focused on the same position on the sample during the whole rotation. The beam spot must always lie on the y-axis of rotation. The sample is rotated from the angle of incidence $70\text{\,}\mathrm{\SIUnitSymbolDegree}$ to $85\text{\,}\mathrm{\SIUnitSymbolDegree}$ in steps of $0.25\text{\,}\mathrm{\SIUnitSymbolDegree}$. This angle setting verifies proper sample alignment and improves the reliability of the evaluation process. To isolate the SH signal from the fundamental, a short-pass filter SP (FESH0800, Thorlabs) is used to remove the fundamental frequency. A Rochon prism analyzer P2 (RPM10, Thorlabs) is required to determine the direction of SHG polarization. A bandpass filter BP (Thorlabs FBH520-40) centered at 520 nm with FWHM 40 nm suppresses parasitic light and higher harmonics. The SHG signal is focused with a biconvex lens L2 (LA1608-A, Thorlabs) with a focal length of 100 mm on the camera (CS165MU1/M, Zelux). The camera calibration factor for calculating the SHG’s power is determined separately by using a Beta-Barium Borate (BBO) crystal as a reference sample. This samples produces a sufficiently strong SHG signal, which can be measured by a calibrated photodiode (S121C, Thorlabs) placed on the camera place. Subsequently, several optical density filters were introduced in the optical path to lower the power of the SH from the BBO crystal to the same order of magnitude as the one from the sample. Knowing the power output and optical density of the filters, we can correlate the power recorded by the camera with the measured signal. The signal was summed up in the area of images with recorded SHG, subsequently divided by the exposure time, and multiplied by the calibration factor of the camera to determine the optical power.
@@
Figure 2: Graphical schema of the optical setup used for second harmonic measurements of the ABC type heterostructures. From the left: femtosecond laser, HW1: half-wave plate, P1: polarizer, HW2: half-wave plate, L1: plano-convex lens, LP: long pass filter, rotation stage with the sample, SP: short pass filter, L2: biconvex lens, P2: Rochon prism analyzer, BP: bandpass filter, camera.
@@
The standard method for evaluating nonlinear second-order susceptibility $\chi^{\textup{(2)}}$ is the Maker fringe technique Maker . It uses the rotation of the sample while the SHG intensity is collected for each angle. However, it is limited to thicker crystals with a minimum size of tens of micrometers to a few millimeters. The limiting factor led to the development of a new method for measuring second-order susceptibility by Hermans et al. On-determination-SHG ; hermans2019 ; koskinen2018 . This method is well suited for evaluating the $\chi^{(2)}$ values of thin films with thicknesses even less than 100 nm where phase mismatch effects can be neglected. Like the Maker fringe technique, the sample is rotated during SHG measurements using a fixed polarization fundamental frequency, and the obtained SH curve is used to fit the nonlinear coefficients. This method does not consider individual separated layers A, B, and C and each interface as a source of second harmonic, but uses an effective medium approximation. Neglecting its birefringent properties, the ABC heterostructure is represented by an average refractive index of $(n_{0}+n_{e})/2$. The SHG is described in terms of the surface second-order nonlinearity $\chi^{\textup{ABC}}_{\textup{s}}$ of the total heterostructure.
Any second-order nonlinear process is described by the second-order susceptibility tensor $\chi^{\textup{(2)}}$, with 27 elements. However, the number of nonzero and independent elements is much smaller due to symmetry properties. The ABC heterostructure belongs to the symmetry group $C_{\infty v}$, as do other amorphous surfaces or thin achiral films with in-plane isotropy. This means that the nonzero coefficients of the second-order susceptibilities are $\chi^{\textup{(2)}}_{{xxz}}=\chi^{\textup{(2)}}_{{yyz}}=\chi^{\textup{(2)}}_{{zxz}}=\chi^{\textup{(2)}}_{{yzy}}$, $\chi^{\textup{(2)}}_{{zxx}}=\chi^{\textup{(2)}}_{{zyy}}$ and $\chi^{\textup{(2)}}_{{zzz}}$, where the $z$ direction is perpendicular to the layers and the $x$ and $y$ directions are longitudinal to the layers. It is also assumed that these coefficients have only real values and that the measurement takes place outside the resonance region, i.e. absorption is not present. The general description of the electric field components of the SH coming from a surface with $C_{\infty v}$ symmetry is described as:
@@
where p and s represent p- and s-polarized components and $E_{\omega}$ is the incoming electric field amplitude of the fundamental frequency. The coefficients $f$, $g$, and $h$ represent generalized layer properties such as nonlinear susceptibility tensor components, thickness, refractive index, and angle of incidence.
The measurement technique is based on the fact that there are two sources of SH signals, $\chi^{\textup{ABC}}_{\textup{s}}$ at the air-thin film interface at the frontside of the sample and $\chi^{\textup{glass}}_{\textup{s}}$ at the glass-air interface at the backside of the sample. Both SH signals interfere with each other, producing specific angle-dependent fringes for this method. Based on prior knowledge of the magnitude of $\chi^{\textup{glass}}_{\textup{s}}$, which is known from either the literature or from calibration measurements, the magnitude of the unknown nonlinearity $\chi^{\textup{ABC}}_{\textup{s}}$ in the thin layer can be deduced. We can obtain the bulk second-order susceptibility as $\chi^{\textup{ABC}}_{\textup{b}}=\chi^{\textup{ABC}}_{\textup{s}}/d$, where $d$ is the total thickness of the heterostructure. Under these approximations made by Hermans et al., a monochromatic p-polarized plane wave of the fundamental frequency of the amplitude of the electric field $E_{\omega,\textup{p}}$ produces a transmitted p-polarized SH electric field $E_{2\omega,\textup{total}}$ as
@@
The parameters for the fundamental frequency (FF) $\omega$ are written in lower case letters, and for the SH in upper case. The parameters $T_{i,j}$ and $t_{i,j}$ are the Fresnel transmission coefficients for p-polarized light that transmits from medium $i$ to medium $j$. The variables $N_{i}$ and $n_{i}$ are refractive indices medium $i$. The variables $\Theta_{i}$ and $\theta_{i}$ are angles of propagation in each medium $i$ with respect to the surface normal $z$. The speed of light is $c$ and the thickness of the substrate is $L_{\textup{glass}}$. The refractive indices for the glass substrate made of pure fused silica are $N_{\textup{glass}}=1.4615$ and $n_{\textup{glass}}=1.4500$. Approximation in Eq. 4 is valid only for materials with low dispersion with $\Theta_{i}\approx\theta_{i}$. Due to this approximation, the values $\chi_{\textup{s},zxx}^{\textup{ABC}}$ and $\chi_{\textup{s},xxz}^{\textup{ABC}}$ cannot be fitted independently, but only as a sum $A_{\textup{s},zx}^{\textup{ABC}}=\chi_{\textup{s},zxx}^{\textup{ABC}}+2\chi_{\textup{s},xxz}^{\textup{ABC}}$. For substrate nonlinearity we used literature values Kauranen_glass_calibration at wavelength 1064 nm $\chi_{\textup{s},zxx}^{\textup{glass}}=3.8\times 10^{-22}$ pm/V, $\chi_{\textup{s},xxz}^{\textup{glass}}=7.9\times 10^{-22}$ pm/V, $\chi_{\textup{s},zzz}^{\textup{glass}}=5.9\times 10^{-22}$ pm/V and we apply the Miller’s rule

### v4

# Systematic Study of Amorphous ABC Heterostructures at the Atomic Scale as a Second-Order Nonlinear Optical Metamaterial

@@
Systematic Study of Amorphous ABC Heterostructures at the Atomic Scale as a Second-Order Nonlinear Optical Metamaterial
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
@@
Systematic exploration of amorphous ABC heterostructures revealed that nanoscale morphological modifications markedly improved their artificial bulk second-order susceptibility. These amorphous birefringent heterostructures were fabricated through plasma-enhanced atomic layer deposition of three oxides, effectively breaking the centrosymmetry on the nanoscale. We observe a dependence of the optical nonlinearity on the thickness variation of three constituent materials, SiO2, TiO2, and Al2O3, ranging from tens of nanometers to the atomic scale, and these thin films exhibit second-order susceptibility at their interfaces. Our findings reveal that the enhancement of nonlinear optical properties is strongly correlated with a high density of layers and superior interface quality, where the interface second-order nonlinearity transitions to bulk-like second-harmonic generation. An effective bulk second-order susceptibility of $\chi_{zzz}=2.0\pm 0.2$ pm/V at the wavelength of 1032 nm is achieved, comparable to typical values for conventional monocrystalline nonlinear materials.
1 Introduction
@@
The rise of atomic layer deposition (ALD) has revolutionized numerous technologies, including transistors [44], memory devices [25], solar cells [40], catalysts [26], batteries [22], and conformal optical coatings [28]. However, its impact on the field of nonlinear optics remains largely untapped [2, 7]. Here, we investigate the potential of ALD-deposited amorphous heterostructures, specifically ABC-type nanolaminates consisting of three distinct dielectrics SiO2 (A), TiO2 (B), and Al2O3 (C) as shown in Fig. 1, to create CMOS-compatible second-order nonlinear metamaterials. The core concept of metamaterials is to engineer materials with properties that are not found in nature by designing their structure at a scale smaller than the wavelength of light they interact with. Atomic layer deposition enables unparalleled precision in the fabrication of the thin films, building one atomic layer at a time with exceptional control and uniformity. Previous studies on layered composites with nonzero bulk second-order susceptibility assumed that only non-centrosymmetric constituents contribute to nonlinearity. However, they neglected interface contributions to the nonlinear susceptibility of the heterostructure [6]. Our goal is to leverage the often overlooked interface, or surface, second-order nonlinearity [13, 36] by stacking three thin oxide layers in 2D heterostructures, thus significantly enhancing the effective nonlinear response by breaking the centrosymmetry as discussed by L. Alloatti et al. and S. Clemmen at al. [2, 7]. Impressively, these heterostructures exhibit a robust, bulk-like nonlinear response, even though all the constituent oxides lack inherent bulk second-order nonlinearity and are amorphous. This produces a highly distinctive material where second-order nonlinear processes arise not from bulk nonlinearity but from surface nonlinearity at the interfaces between centrosymmetric constituents.
Previous studies on ternary oxide heterostructures (Al2O3/TiO2 and SiO2/HfO2) have demonstrated the superior structural and optical properties of ALD-grown thin films [27, 1]. These structures offer temperature and chemical stability, low-temperature growth requirements of $100\penalty 10000\ ^{\circ}\text{C}$, and conformal deposition on a wide range of nano- and micro-structured substrates, without the need to account for substrate compatibility or crystal growth orientation, unlike nonlinear III–V semiconductors such as GaAs [19]. Moreover, they bypass the high-temperature manufacturing requirements of lithium niobate on insulator (LNOI), another widely used nonlinear platform [29]. They also provide the flexibility to engineer the effective refractive index for specific applications by adjusting the ratio of the constituent oxides [27]. Oxide heterostructures exhibit outstanding optical transparency across the visible spectrum to UV spectral range, surpassing the performance of typical III-V technologies. Moreover, the ALD deposition technique and dielectrics used are standard in the semiconductor industry. Altogether, these advantages make ABC heterostructures a powerful platform for integrating photonic structures on a chip using only the commonly available oxides and established semiconductor processes.
@@
Figure 1: The visualization of the ABC-type heterostructures with an optical axis perpendicular to the layers. Changing the density of layers influences their nonlinear optical properties.
@@
In bulk media with inversion symmetry, second-order nonlinear processes are forbidden under the electric dipole approximation, leading to zero bulk second-order susceptibility, $\chi_{\text{bulk}}^{(2)}=0$ [37, 5], and no second harmonic enhancement is observed with varying thickness in pure samples of centrosymmetric materials A, B, or C. However, at the interface between two centrosymmetric media, the inversion symmetry, defined by the transformation $(x,y,z)\to(-x,-y,-z)$, is broken. This symmetry breaking results in a nonzero surface second-order susceptibility, $\chi_{\text{s}}^{(2)}\neq 0$, enabling electric-dipole-allowed second-harmonic generation (SHG) at surfaces and interfaces [35, 11]. ABC heterostructures make use of this principle by stacking nanometer-thick amorphous layers of three centrosymmetric materials A, B, and C. The inversion symmetry is broken locally at each AB, BC, and CA interface. In the case of AB stacks, the centrosymmetry is broken only locally and not globally. However, in ABC stacks, the inversion symmetry is broken both locally at the interfaces and globally, prohibiting total destructive interference of the second harmonic, resulting in a net $\chi_{\text{s}}^{\text{ABC}}$ for the ABC stack. The literature suggests that ABCD stacks, such as those incorporating SiO2/TiO2/Al2O3 with the addition of HfO2, may outperform ABC stacks. However, this potential enhancement remains experimentally unverified [2, 7]. Unfortunately, clear studies are lacking on which combination of inversion-symmetric materials offers the best nonlinear responses. A promising approach is to select materials with high dielectric contrast [13]. However, experimental values for surface second-order susceptibility of materials and interfaces are scarce [32, 41].
@@
Surface nonlinearity arises from three main physical origins. First, the electric dipole contribution due to symmetry breaking at interfaces. Second, a field discontinuity at the interface, where the normal component of the electric field ($E_{z}$) changes rapidly, contributes to the surface nonlinearity via the electric quadrupole (nonlocal) term. Third, structural disparities between materials, such as liquid/solid interfaces, contribute to bulk-like surface nonlinearity even when dielectric constants are matched. The second and third contributions are known as electric quadrupole or nonlocal contributions to surface nonlinear susceptibility [36, 42, 18].
@@
In this work, we present experimental investigations of ABC heterostructures in the form of SiO2/TiO2/Al2O3, deposited by atomic layer deposition. We systematically vary the thickness of each constituent material to evaluate its impact on the efficiency of second-harmonic generation. The structural properties are studied by X-ray reflectivity, scanning transmission electron microscope, and time-of-flight secondary ion mass spectrometry. The film thickness and linear optical properties are determined by spectroscopic ellipsometry. Detailed study of second harmonic generation is carried out to evaluate the second-order nonlinear susceptibility for each morphological composition.
2 Deposition and Characterization
@@
Plasma-enhanced atomic layer deposition (PEALD) was performed using a SILAYO-ICP330 system (Sentech Instruments GmbH) to fabricate SiO2/TiO2/Al2O3 heterostructures. The deposition utilized oxygen plasma as a co-reactant with the following precursors: bis(diethylamino)silane (BDEAS) for SiO2, tetrakis(dimethylamino)titanium (TDMAT) for TiO2, and trimethylaluminum (TMA) for Al2O3. The inductively coupled plasma (ICP) source was operated at $100\penalty 10000\ \text{W}$ for all processes. The reactor chamber was maintained at a temperature of $100\penalty 10000\ ^{\circ}\text{C}$. A 15-minute stabilization period was implemented after sample loading prior to deposition to ensure thermal stability and uniformity. Nitrogen (N2) served as the precursor carrier and purge gas. The growth per cycle (GPC) is determined for single-layer depositions of approximately 20-50 nm. The deposition parameters are presented in Table 1.
@@
[…]
Secondary Ion Mass Spectrometry
[…]
@@
The ordinary and extraordinary refractive indices were determined for each structure using an SE850 DUV variable angle spectroscopic ellipsometer (SENTECH Instruments GmbH), and the data were evaluated in the wavelength range from 400 to 1040nm1040\ $\text${nm} with a uniaxial Sellmeier dispersion model.
@@
The following optical setup was used to investigate the nonlinear optical response of the ABC layers. A femtosecond laser (PHAROS-SP, Light Conversion) emits pulses with full-width half-maxima (FWHM) of $190(10)\text{\,}\mathrm{fs}$ with a central wavelength of 1032 nm and with an average power of up to $1.5\text{\,}\mathrm{W}$ at a repetition rate of $200\text{\,}\mathrm{kHz}$, see Fig. 2. The polarization dependence of the SHG was performed by another femtosecond laser (Satsuma, Amplitude) with a full-width half maxima of $270\text{\,}\mathrm{fs}$, wavelength of 1031.4 nm and repetition rate of $1\text{\,}\mathrm{MHz}$. Power and polarization were controlled using a half-wave plate HW1 (WPH05M-1030, Thorlabs), polarizer P1 (GT10-B, Thorlabs) and a half-wave plate HW2 (Thorlabs AHWP05M-950). The laser beam was focused on the sample using the plano-convex lens L1 (LA1172-B-ML, Thorlabs) with a focal length of 400 mm, producing a beam waist of $w_{0}=55(5)\ \upmu\text{m}$. This leads to a confocal parameter of 18 mm, that is, $2\timesRayleigh$ length. The loosely focused beam allows the front and back sides of the sample to remain in focus during tilting; the sample is 1 mm thick. A longpass filter LP (FELH0900, Thorlabs) was placed behind lens L1 to remove any parasitic light, especially any SHG created by the preceding optical elements. The sample was mounted on an automated rotation stage and aligned so that the laser beam remained focused on the same spot when the sample was tilted. The optical angle of incidence changed from $70\text{\,}\mathrm{\SIUnitSymbolDegree}$ to $85\text{\,}\mathrm{\SIUnitSymbolDegree}$ in steps of $0.25\text{\,}\mathrm{\SIUnitSymbolDegree}$. The generated SH signal from the sample was separated from the fundamental using a shortpass filter SP (FESH0800, Thorlabs). For detecting the generated SH, the sample was imaged with a CMOS camera (CS165MU1/M, Thorlabs) using a biconvex lens with a focal length of 100 mm. Additionally, a band-pass filter (FBH520-40, Thorlabs) was used to suppress the environmental light. A Rochon prism analyzer P2 (RPM10, Thorlabs) was used to determine the direction of SHG polarization. The SH signal was focused using a biconvex lens L2 (LA1608-A, Thorlabs) with a focal length of 100 mm on the camera (CS165MU1/M, Zelux). The camera calibration factor to calculate the power of the SHG is determined separately using a beta-barium borate (BBO) crystal [3].
@@
Figure 2: Graphical schema of the optical setup used for second harmonic measurements of the ABC-type heterostructures. From the left: femtosecond laser, HW1: half-wave plate, P1: polarizer, HW2: half-wave plate, L1: plano-convex lens, LP: longpass filter, tilting stage with the sample, SP: shortpass filter, L2: biconvex lens, P2: Rochon prism analyzer, BP: bandpass filter, camera.
@@
The standard method for evaluating nonlinear second-order susceptibility $\chi^{\textup{(2)}}$ is the Maker fringe technique [20]. It uses the tilting of the sample and the SHG intensity is collected for each angle. However, it is limited to thicker crystals with a minimum size of tens of micrometers to a few millimeters. This limiting factor led to the development of a new method for measuring second-order susceptibility by Hermans et al. [15, 14, 17]. This method is well suited for evaluating the $\chi^{(2)}$ values of thin films with thicknesses even less than 100 nm, where phase mismatch effects can be neglected. Similarly to the Maker fringe technique, the sample was tilted during SHG measurements using a fixed polarization fundamental frequency, and the obtained SH curve was used to fit the nonlinear coefficients. This method does not consider individual separated layers A, B, and C and each interface as a source of second harmonic, but uses an effective medium approximation. Neglecting its birefringent properties, the ABC heterostructure was represented by an average refractive index of $(n_{\textup{o}}+n_{\textup{e}})/2$. SHG was described in terms of the surface second-order nonlinearity $\chi^{\textup{ABC}}_{\textup{s}}$ of the total heterostructure.
Any second-order nonlinear process is described by the second-order susceptibility tensor $\chi^{\textup{(2)}}$, with 27 elements. However, the number of nonzero and independent elements is much smaller due to symmetry properties. The ABC heterostructure belongs to the symmetry group $C_{\infty v}$, as do other amorphous surfaces or thin achiral films with in-plane isotropy. This means that the nonzero coefficients of the second-order susceptibilities are $\chi^{\textup{(2)}}_{{xxz}}=\chi^{\textup{(2)}}_{{yyz}}=\chi^{\textup{(2)}}_{{zxz}}=\chi^{\textup{(2)}}_{{yzy}}$, $\chi^{\textup{(2)}}_{{zxx}}=\chi^{\textup{(2)}}_{{zyy}}$ and $\chi^{\textup{(2)}}_{{zzz}}$, where the $z$ direction is perpendicular to the layers and the $x$ and $y$ directions are parallel to the layers. It is also assumed that these coefficients have only real values and that the measurement takes place outside the resonance region, i.e. absorption is not present. The general description [15] of the electric field components of the SH coming from a surface with $C_{\infty v}$ symmetry is described as:
@@
where p and s represent the p- and s-polarized components and $E_{\omega}$ is the incoming electric field amplitude of the fundamental frequency. The coefficients $f$, $g$, and $h$ represent generalized layer properties such as nonlinear susceptibility tensor components, thickness, refractive index, and angle of incidence.
The measurement technique is based on the fact that there are two sources of SH signals, $\chi^{\textup{ABC}}_{\textup{s}}$ at the air-thin film interface at the frontside of the sample and $\chi^{\textup{glass}}_{\textup{s}}$ at the glass-air interface at the backside of the sample. Both SH signals interfere with each other, producing specific angle-dependent fringes for this method. Based on prior knowledge of the magnitude of $\chi^{\textup{glass}}_{\textup{s}}$, which is known from either the literature or from calibration measurements, the magnitude of the unknown nonlinearity $\chi^{\textup{ABC}}_{\textup{s}}$ in the thin layer can be deduced. We can obtain the bulk second-order susceptibility as $\chi^{\textup{ABC}}_{\textup{b}}=\chi^{\textup{ABC}}_{\textup{s}}/d$, where $d$ is the total thickness of the heterostructure. Under these approximations made by Hermans et al.[15, 14, 17], a monochromatic p-polarized plane wave of the fundamental frequency of the amplitude of the electric field $E_{\omega,\textup{p}}$ produces a transmitted p-polarized SH electric field $E_{2\omega,\textup{total}}$ as
@@
The parameters for the fundamental frequency (FF) $\omega$ are denoted by lower case letters, and for the SH by upper case. The parameters $T_{i,j}$ and $t_{i,j}$ are the Fresnel transmission coefficients for p-polarized light that transmits from medium $i$ to medium $j$. The variables $N_{i}$ and $n_{i}$ are refractive indices of medium $i$. The variables $\Theta_{i}$ and $\theta_{i}$ are angles of propagation in each medium $i$ with respect to the surface normal $z$. The speed of light is $c$ and the thickness of the substrate is $L_{\textup{glass}}$. The refractive indices for the glass substrate made of pure fused silica are $N_{\textup{glass}}=1.4615$ and $n_{\textup{glass}}=1.4500$. The approximation in Eq. 4 is valid only for materials with low dispersion with $\Theta_{i}\approx\theta_{i}$. Due to this approximation, the values $\chi_{\textup{s},zxx}^{\textup{ABC}}$ and $\chi_{\textup{s},xxz}^{\textup{ABC}}$ cannot be fitted independently, but only as a sum $A_{\textup{s},zx}^{\textup{ABC}}=\chi_{\textup{s},zxx}^{\textup{ABC}}+2\chi_{\textup{s},xxz}^{\textup{ABC}}$. For substrate nonlinearity we used literature values [32] at the wavelength 1064 nm: $\chi_{\textup{s},zxx}^{\textup{glass}}=3.8\times 10^{-22}$ pm/V, $\chi_{\textup{s},xxz}^{\textup{glass}}=7.9\times 10^{-22}$ pm/V, $\chi_{\textup{s},zzz}^{\textup{glass}}=5.9\times 10^{-22}$ pm/V and we apply the Miller’s rule